OpenStudy (anonymous):

Calculate the double integral where R = {(x, y) ∈ R 2 | 0 ≤ x ≤ 1, −3 ≤ y ≤ 3 }.

OpenStudy (anonymous):

\[\int\limits_{R}^{}\frac{ xy^2 }{ x^2+1 }d \lambda2\]

OpenStudy (anonymous):

Does it become this, but interpreting the domains:

OpenStudy (anonymous):

\[\int\limits_{0}^{1}(\int\limits_{-3}^{3}\frac{ xy^2 }{ x^2+1 }dy)dx\]